# Learning to generate feasible graphs using graph grammars

### v1

@@ Abstract
Generative methods for graphs need to be sufficiently flexible to model complex dependencies between sets of nodes. At the same time, the generated graphs need to satisfy domain-dependent feasibility conditions, that is, they should not violate certain constraints that would make their interpretation impossible within the given application domain (e.g. a molecular graph where an atom has a very large number of chemical bounds). Crucially, constraints can involve not only local but also long-range dependencies: for example, the maximal length of a cycle can be bounded.
@@ 1 Introduction
In Costa (2017) the problem of generating elements of a structured domain was framed as the equivalent problem of sampling from a corresponding underlying probability distribution defined over a (learned) class of structures. Specifically, they employ a context-sensitive grammar to accurately model complex dependencies between different parts of an instance. The authors acknowledge that an approach based exclusively on a grammar is not sufficient since the number of proposed graphs grows exponentially with the number of production rules in the grammar. To address the issue, it was proposed to use a Metropolis Hastings (MH) Markov Chain Monte Carlo (MCMC) method, where the sampling problem is reduced to the easier task of simulation. They use context-sensitive graph grammar to inform the MH proposal distribution, but also introduce a probability density estimator to define the MH acceptance procedure. This allows us to deal separately with local and global constraints: the locally context-sensitive graph grammar is used for the local constraints and the regularized statistical model is used for the global or long range constraints. The two approaches complement each other: the grammar is a flexible non-parametric approach that can model complex dependencies between vertices that are within a short path distance from each other; the statistical model, instead, can employ the bias derived from the particular functional family (linear) and the type of regularization used (a penalty over the $\ell_{2}$ norm of the parameters) to generalize long range dependencies to similar cases. This approach is therefore adequate when the underlying concept exhibit local dependencies that are more complex than long range ones. Unfortunately in some application domains instances can exhibit complex long range dependencies between their different constituent parts. This is the case for RNA polymers, long sequences of atomic entities (nucleotides) that self interact, establishing pairwise bounds that can typically span the entire length of the sequence, where sequences can range from hundreds to thousands of atomic parts.
@@ 2.1 Contribution
Figure 1: A depiction of core and interface subgraphs. Dark shading indicates the core subgraph, while light shading represents the interface. The diagram on the left shows a molecular graph with its core substitution, while the middle and right diagrams depict RNA encodings at different resolutions (nucleotide-level and structural-element-level, respectively). At nucleotide-level nodes are labeled with nucleotides codes ACGT, while at structural level nodes are labeled as H)airpin, M)ultiloop, S)tem, D)angling end.
Constraint integration. It should be possible to easily make use of specific feasibility constraints when they are available for a given domain. For RNAs the relation between sequence and structure is known to be governed by thermodynamical forces that seek to minimize the amount of free energy of the molecule. The structure of a given sequence can be computed using sophisticated dynamic programming optimization algorithms. To integrate these constraints in the constructive protocol we let the user specify a transformation function which maps candidate instances to feasible instances. In our case, the transformation takes a graph representing an RNA structure constructed by the proposed procedure, removes the pairwise bounds, and recomputes them as the solution of a given folding algorithm. This transformation ensures that we are always computing the acceptance probability on viable RNA structures.
@@ 3 Empirical evaluation
Figure 4: Estimated equivalence by Infernal. The horizontal line is the threshold for biological significance.
Figure 5: Similarity to the training instances. The generated graphs should differ from the training set, yet induce the same probability density.
@@ 3.1 Graph generative approaches
We examine the drawbacks of each approach. VAE and ORGAN produce less than 10% valid graphs, CharRNN and AAE produce 35% and 36% valid and unique graphs. The best performing approach is CharNN with top performance in FCD, Frag and Scaf measures. The drawback of CharNN is that it produces molecules of such small size (5-7 atoms) that exhaustive enumeration becomes possible. AAE performs best in SNN (this is probably due generating graphs of the appropriate size), but low according to the IntDiv scores(0.69/0.66, while all other generators are in the range 0.84 $\pm$ 0.02), suggesting that its products are very similar to the training material (it outputs only 216 unique graphs over 604 valid ones). The products of the non coarsened grammar perform very similar to the coarsened version however only 79 of the 1000 generated graphs pass the filter test (second lowest is ORGAN with 58.3% , coarsened grammar: 87.2%) this suggests long range dependencies are not met, resulting in unstable molecules. Each weakness of our approach is that the generated instances tend to be large in size which affects the SNN measure. This tendency is a byproduct of the acceptance procedure and in future work will be addressed. In summary, our procedure, while being domain agnostic, generates 100% viable graphs, with comparable performance to the other specialized algorithms (see Tab. 1).
@@
Table 1: FCD - ChemNet distance; a measurement for molecule likeness, Frag/Scaf - cosine similarity to graph fragments, SNN - similarity to the graphs in the test set, IntDiv - diversity of generated artifacts, Filters - fraction of graphs that is non toxic or not unstable. The graph coarsened graph grammar approach performs similar to the specialized methods while generating only valid graphs. CharRNN generates very small graphs, AAE generates graphs that are very similar to those in the training set.
GraphRNN (You et al., 2018) works with unlabeled graphs and sets out to model long range dependencies using an auto regressive model. We Trained this method on the seed sequences of RF01725 whose structure we obtained with the MUSCLE scheme described previously. In Fig. 6 we show three graphs generated by GraphRNN. While the hairpin structure is present, in the lower part of the figure we see cycles and in general more than two dangling ends indicating that the most basic constraint, a continuous ribose backbone, is violated. Out of 1024 generated graphs only 13 passed the following simple filters: a) maximal 2 nodes have a degree of 1. b) none has a degree > 3. c) is connected. d) cycles where exactly 2 nodes have degree 3 and any other vertex has degree 2, the degree 3 nodes must be directly connected. a) and c) are necessary for RNA graphs because an nucleotides form a path graph on top of which H-bonds (edges between non adjacent vertices) are formed. One nucleotide can only form a single H-bond (filter b). Filter d eliminates "nearly-hairpin" structures as seen in Fig. 6(lower left), where the degree 3 nodes next to each other, it would be a valid hairpin structure.

### v2

@@ Abstract
Generative methods for graphs need to be sufficiently flexible to model complex dependencies between sets of nodes. At the same time, the generated graphs need to satisfy domain-dependent feasibility conditions, that is, they should not violate certain constraints that would make their interpretation impossible within the given application domain (e.g. a molecular graph where an atom has a very large number of chemical bonds). Crucially, constraints can involve not only local but also long-range dependencies: for example, the maximal length of a cycle can be bounded.
@@ 1 Introduction
In Costa (2017) the problem of generating elements of a structured domain was framed as the equivalent problem of sampling from a corresponding underlying probability distribution defined over a (learned) class of structures. Specifically, they employ a context-sensitive grammar to accurately model complex dependencies between different parts of an instance. The authors acknowledge that an approach based exclusively on a grammar is not sufficient since the number of proposed graphs grows exponentially with the number of production rules in the grammar. To address the issue, it was proposed to use a Metropolis Hastings (MH) Markov Chain Monte Carlo (MCMC) method, where the sampling problem is reduced to the easier task of simulation. They use context-sensitive graph grammar to inform the MH proposal distribution, but also introduce a probability density estimator to define the MH acceptance procedure. This allows us to deal separately with local and global constraints: the locally context-sensitive graph grammar is used for the local constraints and the regularized statistical model is used for the global or long range constraints. The two approaches complement each other: the grammar is a flexible non-parametric approach that can model complex dependencies between vertices that are within a short path distance from each other; the statistical model, instead, can employ the bias derived from the particular functional family (linear) and the type of regularization used (a penalty over the $\ell_{2}$ norm of the parameters) to generalize long range dependencies to similar cases. This approach is therefore adequate when the underlying concept exhibit local dependencies that are more complex than long range ones. Unfortunately in some application domains instances can exhibit complex long range dependencies between their different constituent parts. This is the case for RNA polymers, long sequences of atomic entities (nucleotides) that self interact, establishing pairwise hydrogen bonds that can typically span the entire length of the sequence, where sequences can range from hundreds to thousands of atomic parts.
@@ 2.1 Contribution
Figure 1: A depiction of core and interface subgraphs. Dark shading indicates the core subgraph, while light shading represents the interface. The diagram on the left shows a molecular graph with its core substitution, while the middle and right diagrams depict RNA encodings at different resolutions (nucleotide-level and structural-element-level, respectively). At nucleotide-level nodes are labeled with nucleotides codes ACGU, while at structural level nodes are labeled as H)airpin, M)ultiloop, S)tem, D)angling end.
Constraint integration. It should be possible to easily make use of specific feasibility constraints when they are available for a given domain. For RNAs the relation between sequence and structure is known to be governed by thermodynamical forces that seek to minimize the amount of free energy of the molecule. The structure of a given sequence can be computed using sophisticated dynamic programming optimization algorithms. To integrate these constraints in the constructive protocol we let the user specify a transformation function which maps candidate instances to feasible instances. In our case, the transformation takes a graph representing an RNA structure constructed by the proposed procedure, removes the pairwise nucleotide bonds, and recomputes them as the solution of a given folding algorithm. This transformation ensures that we are always computing the acceptance probability on viable RNA structures.
@@ 3 Empirical evaluation
Figure 4: Evaluation of generated RNAs. The Infernal bit score is the expert-models classification score. The horizontal line is the threshold for biological significance associated with this RNA family.
Figure 5: Evaluation of generated RNAs. Here we compare the generated sequences to the training material. The generated graphs should differ from the training set, yet induce the same probability density.
@@ 3.1 Graph generative approaches
We examine the drawbacks of each approach. VAE and ORGAN produce less than 10% valid graphs, CharRNN and AAE produce 35% and 36% valid and unique graphs. The best performing approach is CharNN with top performance in FCD, Frag and Scaf measures. The drawback of CharNN is that it produces molecules of such small size (5-7 atoms) that exhaustive enumeration becomes possible. AAE performs best in SNN (this is likely due to generating graphs of the appropriate size), but low according to the IntDiv scores(0.69/0.66, while all other generators are in the range 0.84 $\pm$ 0.02), suggesting that its products are very similar to the training material (it outputs only 216 unique graphs over 604 valid ones). The products of the non coarsened grammar perform very similar to the coarsened version however only 79 of the 1000 generated graphs pass the filter test (second lowest is ORGAN with 58.3% , coarsened grammar: 87.2%) this suggests long range dependencies are not met, resulting in unstable molecules. The weakness of our approach is that the generated instances tend to be large in size which affects the SNN measure. This tendency is a byproduct of the acceptance procedure and in future work will be addressed. In summary, our procedure, while being domain agnostic, generates 100% viable graphs, with comparable performance to the other specialized algorithms (see Tab. 1).
@@
Table 1:
Results from the MOSES benchmarks comparing various methods for molecule generation. FCD - ChemNet distance; a measurement for molecule likeness, Frag/Scaf - cosine similarity to graph fragments, SNN - similarity to the graphs in the test set, IntDiv - diversity of generated artifacts, Filters - fraction of graphs that is non toxic or not unstable. The graph coarsened graph grammar approach performs similar to the specialized methods while generating only valid graphs. CharRNN generates very small graphs, AAE generates graphs that are very similar to those in the training set.
GraphRNN (You et al., 2018) works with unlabeled graphs and sets out to model long range dependencies using an auto regressive model. We Trained this method on the seed sequences of RF01725 whose structure we obtained with the MUSCLE scheme described previously. In Fig. 6 we show three graphs generated by GraphRNN. While the hairpin structure is present, in the lower part of the figure we see cycles and in general more than two dangling ends indicating that the most basic constraint, a continuous ribose backbone, is violated. Out of 1024 generated graphs only 13 passed the following simple filters: a) maximal 2 nodes have a degree of 1. b) none has a degree > 3. c) is connected. d) cycles where exactly 2 nodes have degree 3 and any other vertex has degree 2, the degree 3 nodes must be directly connected. a) and c) are necessary for RNA graphs because an nucleotides form a path graph on top of which H-bonds (edges between non adjacent vertices) are formed. One nucleotide can only form a single H-bond (filter b). Filter d eliminates "nearly-hairpin" structures as seen in Fig. 6(lower left), were the degree 3 nodes next to each other, it would form a valid hairpin structure.